# Neural Approximate Mirror Maps for Constrained Diffusion Models

### v1

@@ 5 Conclusion
Limitations
Our approach does not guarantee exact constraint satisfaction. The advantage is greater flexibility with the constraint, as we only require a differentiable constraint distance. As we saw with the count constraint, the constraint loss may not always be easy to optimize via gradient-based methods. Finally, there are many fruitful theoretical questions to pursue, such as formulating guarantees on the quality of the learned mirror map in relation to different classes of constraints.
Summary
We have proposed a method for constrained diffusion models that minimally restricts the generative model and constraint. A NAMM consists of a mirror map $\mathbf{g}_{\phi}$ and its approximate inverse $\mathbf{f}_{\psi}$, which is robust to noise added to samples in the mirror space induced by $\mathbf{g}_{\phi}$. One can train a mirror diffusion model in this mirror space to generate samples that are constrained by construction via the inverse map. We have validated that our method provides significantly better constraint satisfaction than a vanilla diffusion model on various constraints and have shown its utility for solving constrained inverse problems. Our work establishes that neural approximate mirror maps effectively rein in generative models according to visually-subtle yet physically-meaningful constraints.
@@ Acknowledgments
The authors would like to thank Rob Webber, Jamie Smith, Dmitrii Kochkov, Alex Ogren, and Florian Schaefer for their helpful discussions. BTF and KLB acknowledge support from the NSF GRFP, NSF Award 2048237, and the Amazon AI4Science Partnership Discovery Grant. RB acknowledges support from the US Department of Energy AEOLUS center (award DE-SC0019303), the Air Force Office of Scientific Research MURI on “Machine Learning and Physics-Based Modeling and Simulation” (award FA9550-20-1-0358), and a Department of Defense (DoD) Vannevar Bush Faculty Fellowship (award N00014-22-1-2790).
Appendix A Experiment details
A.1 Implementation
@@
For training the score model $s_{\theta}$ in the learned mirror space, we followed the implementation of Song et al. [61]. We used the NCSN++ architecture with $64$ filters in the first layer and the VP SDE with $\beta_{\min}=0.1$ and $\beta_{\max}=20$. Training was done using the Adam optimizer with a learning rate of $0.0002$ and gradient norm clipping with a threshold of $1$.
@@
For $\mathbf{g}_{\phi}$, we followed the implementation of the gradient of a strongly-convex ICNN of Tan et al. [65], configuring the ICNN to be $0.9$-strongly convex. Following the settings of CycleGAN [73], $\mathbf{f}_{\psi}$ was implemented as a ResNet-based generator with $6$ residual blocks and $32$ filters in the last convolutional layer. For all constraints except the divergence-free constraint, we had the ResNet-based generator output the residual image (i.e., $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})=\text{ResNet}(\tilde{\mathbf{x}})+\tilde{\mathbf{x}}$). We found that for the divergence-free constraint, a non-residual-based inverse map (i.e., $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})=\text{ResNet}(\tilde{\mathbf{x}})$) achieves better constraint loss. The NAMM was trained using Adam optimizer with a learning rate of $0.001$ for the divergence-free constraint and a learning rate of $0.0002$ for all other demonstrated constraints.
Tab. 2 shows the hyperparameter choices for each constraint. The regularization weight $\lambda_{\text{reg}}$ in the NAMM objective (Eq. 1) was fixed at $0.001$. We used $3$ ICNN layers for images $64\times 64$ or smaller and $2$ ICNN layers for images $128\times 128$ or larger for the sake of efficiency. These hyperparameter values do not need to be heavily tuned, as we chose these settings through a coarse parameter search (e.g., trying $\lambda_{\text{constr}}=0.01$ or $\lambda=1$ to see which would lead to reasonable loss curves).
@@
The main results shown in Fig. 3 were taken from the finetuned NAMM, ensuring that the total training time of the NAMM, MDM, and finetuning did not exceed the total training time of the baseline DM. While we kept track of the validation loss, this was not used to determine stopping time. We found that the NAMM training and MDM training were not prone to overfitting, so we chose the total number of epochs based on observing a reasonable level of convergence of the loss curves. We found that some overfitting is possible during finetuning but did not perform early stopping. All results were obtained from unseen test data because we fed random samples from the MDM into the inverse map and made sure not to use the same random seed as the one used to generate finetuning data. For all constraints, we generated $12800$ training examples from the MDM for finetuning.
Tab. 3 details the exact number of training epochs for each constraint. Fig. 4 in the main results compares the constraint distances of our method without finetuning, our method with finetuning, and the baseline DM as a function of compute time.
@@
Total Flux $30$ $200$ $970$ $450$
1D Burgers’ $100$ $300$ $600$ $1500$
Divergence-free $100$ $300$ $400$ $2000$
Periodic $50$ $300$ $9050$ $1000$
Count $50$ $300$ $650$ $1500$
@@
For the comparison of parameterizing the mirror map as the gradient of an ICNN versus as a ResNet-based generator (Fig. 7), we used a ResNet-based generator that outputs the residual image. This means that the inverse mirror map was parameterized as a residual-based network ($\mathbf{f}_{\psi}(\tilde{\mathbf{x}})=\text{ResNet}(\tilde{\mathbf{x}};\psi)+\tilde{\mathbf{x}}$) and so was the ResNet-based forward mirror map ($\mathbf{g}_{\phi}(\mathbf{x})=\text{ResNet}(\mathbf{x};\phi)+\mathbf{x}$).
A.2 Ablation of sequence of noisy mirror distributions
Recall that the NAMM training objective involves optimizing $\mathbf{f}_{\psi}$ over the sequence of noisy mirror distributions defined in Eq. 2. Thus instead of considering a single noisy mirror distribution ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}\ast\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ with a fixed noise level $\sigma$, we perturb samples from ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$ with varying levels of Gaussian noise with variances ranging from 00 to $\sigma_{\text{max}}^{2}$. In Fig. 8, we compare this choice, which involves integrating over $\sigma\in[0,\sigma_{\text{max}}]$, to the use of a fixed $\sigma=\sigma_{\text{max}}$.
Figure 8: Fixed $\sigma$ vs. integrating over $[0,\sigma_{\text{max}}]$. The NAMM objective involves optimizing $\mathbf{f}_{\psi}$ over the sequence of noisy mirror distributions defined in Eq. 2. We compare this approach of integrating over $\sigma\in[0,\sigma_{\text{max}}]$ to setting a fixed noise std. dev. $\sigma=\sigma_{\text{max}}$ in the context of the 1D Burgers’ constraint (here $\sigma_{\text{max}}=0.1$). Both before and after finetuning, the constraint distances $\ell_{\text{burgers}}$ of inverted MDM samples are smaller if the NAMM was trained with varying $\sigma$. The histograms show the constraint distances of $128$ samples from each method.
A.3 Data assimilation with mirror DPS
Following the original DPS [17], we use a measurement-weight hyperparameter $\zeta\in\mathbb{R}_{>0}$ to re-weight the time-dependent measurement likelihood. At diffusion time $t$, the measurement weight is given by

### v2

@@ 5 Conclusion
We have proposed a method for constrained diffusion models that minimally restricts the generative model and constraint. A NAMM consists of a mirror map $\mathbf{g}_{\phi}$ and its approximate inverse $\mathbf{f}_{\psi}$, which is robust to noise added to samples in the mirror space induced by $\mathbf{g}_{\phi}$. One can train a mirror diffusion model in this mirror space to generate samples that are constrained by construction via the inverse map. We have validated that our method provides significantly better constraint satisfaction than a vanilla diffusion model on physics-based, geometric, and semantic constraints and have shown its utility for solving constrained inverse problems. Our work establishes that NAMMs effectively rein in generative models according to visually-subtle yet physically-meaningful constraints.
@@ Acknowledgments
The authors would like to thank Rob Webber, Jamie Smith, Dmitrii Kochkov, Alex Ogren, and Florian Schäfer for their helpful discussions. BTF and KLB acknowledge support from the NSF GRFP, NSF Award 2048237, NSF Award 2034306, the Pritzker Award, and the Amazon AI4Science Partnership Discovery Grant. RB acknowledges support from the US Department of Energy AEOLUS center (award DE-SC0019303), the Air Force Office of Scientific Research MURI on “Machine Learning and Physics-Based Modeling and Simulation” (award FA9550-20-1-0358), and a Department of Defense (DoD) Vannevar Bush Faculty Fellowship (award N00014-22-1-2790).
Appendix A NAMMs for constrained VAEs
Our approach is compatible with any generative model, not just diffusion models. Once a NAMM is trained, any generative model can be trained in the learned mirror space and its samples mapped back to the constrained space via the learned inverse mirror map. In this appendix, we apply our approach to training a variational autoencoder (VAE) that satisfies the divergence-free constraint, comparing a VAE trained in the learned mirror space (“MVAE”) to a VAE trained in the original data space.
A.1 Improved constraint satisfaction with a mirror VAE
For this experiment, we trained a VAE in the mirror space induced by the learned mirror map that was trained for the divergence-free constraint (without finetuning). We call this the “mirror VAE,” or MVAE, approach. As a baseline, we trained the same VAE architecture on the original divergence-free data without transformation. We note that the same data are used to train both the MVAE and VAE; the only difference is that the MVAE is trained in the mirror space, while the VAE is trained in the original space. The training procedure was otherwise the same for both the MVAE and VAE.
Figure 8 shows samples from the MVAE and VAE. We ensured that the total training time of the MVAE did not exceed that of the VAE (on the same hardware, 4×4$\times$4 × A100 GPUs). The VAE was trained for $3500$ epochs, and the MVAE was trained for $600$ epochs (following $100$ epochs of NAMM training). Both approaches produce visually similar samples, yet the images of the divergence field and histograms for this constraint distance show that the MVAE leads to overall better constraint satisfaction. Furthermore, in terms of MMD and KID, the MVAE distribution is even closer to the true data distribution.
Figure 8: Mirror VAE (MVAE) vs. VAE, demonstrated on the divergence-free constraint. Baseline samples are obtained from a VAE trained in the original space, while MVAE samples are obtained by sampling from a VAE trained in the mirror space and then mapping those samples back to the original space via the learned inverse mirror map. Three samples are shown (vorticity on the left and divergence on the right) for each method. (Recall that each image consists of eight state snapshots; here we have labeled the number of each snapshot.) The vorticity fields show that the visual statistics of both generated distributions are extremely similar, but the corresponding divergence fields are drastically different. The MVAE samples are much closer to satisfying 00-divergence everywhere. As further evidence, the histograms show that normalized constraint distances of MVAE samples are significantly lower. We also report the mean $\pm$ std. dev. constraint distance (CD), computed as $100\lambda_{\text{constr}}\ell$, as well as the MMD and KID. All three metrics were estimated with $10000$ generated and true samples. The MVAE leads to improved constraint satisfaction and distribution-matching accuracy compared to a vanilla VAE. This experiment demonstrates how a NAMM can be used to constrain generative models besides diffusion models.
A.2 VAE implementation details
We used a convolutional autoencoder architecture consisting of five convolutional layers with GELU activation functions in the encoder and decoder. Our implementation is borrowed from the autoencoder tutorial of Lippe (2024). We set the number of latent dimensions to $128$ and the number of features in the first layer of the encoder to $64$. For training, we followed the $\beta$-VAE training objective (Higgins et al., 2017), which consists of two terms: one to increase the likelihood of training data under the VAE probabilistic model and one to minimize the KL divergence from the latent distribution to a Gaussian prior. The latter is weighted by a scalar $\beta>0$. For our purposes, the maximum-likelihood term corresponds to a mean-squared-error (MSE) reconstruction loss, and we set $\beta=0.001$. We used the Adam optimizer with a learning rate of $0.0002$.
Appendix B Sample comparisons
Figure 9 shows random samples from our approach and the baseline DM approach, corresponding to the results shown in Figure 3. We show more samples in this appendix to give a sense of the visual similarity between the samples generated with our approach and those generated with the baseline approach. We emphasize again that our samples are much closer to the constraint set despite being visually indistinguishable from baseline samples.
Figure 9: Comparisons of generated distributions. Nine (six for the divergence-free constraint) samples are shown for each distribution. The constraint distance (CD), MMD, and KID values are repeated here from Tab. 1 for ease of comparison. Qualitatively, our approach gives samples that are visually very similar to real samples and baseline samples.
Appendix C Constraint hyperparameters
Here we provide an ablation study of the constraint hyperparameters, $\lambda_{\text{constr}}$ and $\sigma_{\text{max}}$, for all the demonstrated constraints. We performed a hyperparameter sweep across all combinations of $\lambda_{\text{constr}}\in[0.01,0.1,1.0]$ and $\sigma_{\text{max}}\in[0.001,0.1,0.5]$. For each setting, we trained a NAMM and then an MDM, and then we evaluated the constraint satisfaction and distribution-matching accuracy of the MDM samples. Table 2 reports the constraint distance, MMD, and KID metrics for all hyperparameter settings and constraints.
These results illustrate how performance changes with respect to $\lambda_{\text{constr}}$ and $\sigma_{\text{max}}$. We observe that increasing $\lambda_{\text{constr}}$ generally leads to lower constraint distances for any constraint. For some constraints, there is also an improvement in MMD/KID as $\lambda_{\text{constr}}$ increases (e.g., 1D Burgers’), whereas in other cases, there seems to be a tradeoff between constraint satisfaction and distribution-matching accuracy (e.g., Count). We observe that the metrics change non-monotonically as a function of $\sigma_{\text{max}}$. This may be attributed to the fact that the training objective has a nonlinear dependence on $\sigma_{\text{max}}$. When applying our method to a new constraint, one can run a similar hyperparameter sweep to choose values that lead to the best combination of constraint satisfaction and distribution-matching accuracy.
$\sigma_{\text{max}}=0.001$ $\sigma_{\text{max}}=0.1$ $\sigma_{\text{max}}=0.5$
Total Brightness $\lambda_{\text{constr}}=0.01$ CD ($\downarrow$) $0.66\pm 0.38$ $0.55\pm 0.45$ $0.66\pm 0.57$
MMD ($\downarrow$) $0.0517\pm 0.0060$ $0.0549\pm 0.0067$ $0.0757\pm 0.0055$
KID ($\downarrow$) $0.0386\pm 0.0025$ $0.0078\pm 0.0014$ $0.0062\pm 0.0008$
$\lambda_{\text{constr}}=0.1$ CD ($\downarrow$) $0.03\pm 0.02$ $0.02\pm 0.01$ $0.22\pm 0.03$
MMD ($\downarrow$) $1.1723\pm 0.0006$ $1.1723\pm 0.0006$ $1.1727\pm 0.0006$
KID ($\downarrow$) $0.4372\pm 0.0043$ $0.5275\pm 0.0045$ $0.6396\pm 0.0047$
$\lambda_{\text{constr}}=1.0$ CD ($\downarrow$) $0.07\pm 0.02$ $0.03\pm 0.01$ $0.11\pm 0.01$
MMD ($\downarrow$) $1.1721\pm 0.0006$ $1.1722\pm 0.0006$ $1.1727\pm 0.0006$
KID ($\downarrow$) $0.5052\pm 0.0044$ $0.5191\pm 0.0045$ $0.6151\pm 0.0046$
1D Burgers’ $\lambda_{\text{constr}}=0.01$ CD ($\downarrow$) $0.20\pm 0.03$ $0.27\pm 0.03$ $0.26\pm 0.03$
MMD ($\downarrow$) $0.0488\pm 0.0090$ $0.0631\pm 0.0095$ $0.0945\pm 0.0078$
KID ($\downarrow$) $0.0483\pm 0.0012$ $0.1111\pm 0.0018$ $0.1142\pm 0.0020$
$\lambda_{\text{constr}}=0.1$ CD ($\downarrow$) $0.20\pm 0.03$ $0.20\pm 0.02$ $0.18\pm 0.02$
MMD ($\downarrow$) $0.0784\pm 0.0075$ $0.0682\pm 0.0071$ $0.1069\pm 0.0062$
KID ($\downarrow$) $0.0434\pm 0.0014$ $0.0495\pm 0.0016$ $0.0499\pm 0.0015$
$\lambda_{\text{constr}}=1.0$ CD ($\downarrow$) $0.20\pm 0.03$ $0.07\pm 0.01$ $0.17\pm 0.02$
MMD ($\downarrow$) $0.1159\pm 0.0078$ $0.1092\pm 0.0090$ $0.2169\pm 0.0084$
KID ($\downarrow$) $0.0494\pm 0.0012$ $0.0026\pm 0.0004$ $0.0144\pm 0.0008$
Divergence-free $\lambda_{\text{constr}}=0.01$ CD ($\downarrow$) $4.17\pm 0.29$ $5.17\pm 0.33$ $6.01\pm 0.35$
MMD ($\downarrow$) $0.0493\pm 0.0029$ $0.0484\pm 0.0025$ $0.0522\pm 0.0018$
KID ($\downarrow$) $0.0007\pm 0.0002$ $0.0029\pm 0.0004$ $0.0062\pm 0.0008$
$\lambda_{\text{constr}}=0.1$ CD ($\downarrow$) $3.37\pm 0.30$ $3.16\pm 0.33$ $3.56\pm 0.31$
MMD ($\downarrow$) $0.0593\pm 0.0026$ $0.0569\pm 0.0031$ $0.0588\pm 0.0030$
KID ($\downarrow$) $0.0029\pm 0.0006$ $0.0009\pm 0.0002$ $0.0020\pm 0.0003$
$\lambda_{\text{constr}}=1.0$ CD ($\downarrow$) $2.42\pm 0.36$ $2.28\pm 0.18$ $2.61\pm 0.21$
MMD ($\downarrow$) $0.0554\pm 0.0037$ $0.0548\pm 0.0031$ $0.0593\pm 0.0019$
KID ($\downarrow$) $0.0031\pm 0.0005$ $0.0031\pm 0.0004$ $0.0029\pm 0.0004$
Periodic $\lambda_{\text{constr}}=0.01$ CD ($\downarrow$) $0.25\pm 0.40$ $0.28\pm 0.12$ $0.30\pm 0.15$
MMD ($\downarrow$) $0.2133\pm 0.0057$ $0.0484\pm 0.0041$ $0.0507\pm 0.0039$
KID ($\downarrow$) $0.0215\pm 0.0018$ $0.0006\pm 0.0002$ $0.0019\pm 0.0004$
$\lambda_{\text{constr}}=0.1$ CD ($\downarrow$) $0.35\pm 0.29$ $0.15\pm 0.09$ $0.35\pm 0.20$
MMD ($\downarrow$) $0.0795\pm 0.0059$ $0.0621\pm 0.0045$ $0.1013\pm 0.0060$
KID ($\downarrow$) $0.0054\pm 0.0008$ $0.0015\pm 0.0005$ $0.0044\pm 0.0008$
$\lambda_{\text{constr}}=1.0$ CD ($\downarrow$) $0.30\pm 0.11$ $0.05\pm 0.06$ $00.00\pm 0$ 0.00
MMD ($\downarrow$) $0.1092\pm 0.0056$ $0.0856\pm 0.0057$ $0.7822\pm 0.0037$
KID ($\downarrow$) $0.0049\pm 0.0007$ $0.0044\pm 0.0009$ $0.8408\pm 0.0020$
Count $\lambda_{\text{constr}}=0.01$ CD ($\downarrow$) $0.73\pm 0.55$ $0.65\pm 0.49$ $0.61\pm 0.47$
MMD ($\downarrow$) $0.0715\pm 0.0057$ $0.0449\pm 0.0030$ $0.2119\pm 0.0067$
KID ($\downarrow$) $0.0084\pm 0.0006$ $0.0514\pm 0.0011$ $0.0266\pm 0.0011$
$\lambda_{\text{constr}}=0.1$ CD ($\downarrow$) $0.13\pm 0.10$ $0.59\pm 0.45$ $0.52\pm 0.41$
MMD ($\downarrow$) $0.1773\pm 0.0023$ $0.0884\pm 0.0058$ $0.0520\pm 0.0032$
KID ($\downarrow$) $0.5006\pm 0.0040$ $0.0594\pm 0.0012$ $0.0086\pm 0.0004$
$\lambda_{\text{constr}}=1.0$ CD ($\downarrow$) $0.01\pm 0.01$ $0.02\pm 0.01$ $0.03\pm 0.01$
MMD ($\downarrow$) $0.6346\pm 0.0022$ $0.6774\pm 0.0024$ $0.7988\pm 0.0029$
KID ($\downarrow$) $0.4995\pm 0.0018$ $0.4957\pm 0.0020$ $0.4786\pm 0.0016$
Table 2: Ablation study of constraint loss hyperparameters. Constr. dist. (CD) $=100\lambda_{\text{constr}}\ell$. MMD and KID are metrics for distribution-matching accuracy. The mean $\pm$ std. dev. of each metric was estimated with $10000$ samples. For each constraint, the NAMM and MDM were trained for as many epochs (before finetuning) as reported in Table 4. The hyperparameter setting used for the main paper results reported in Table 1 is underlined for each constraint (the results here are slightly different due to randomness in the training runs). These results, which were attained without finetuning, give a sense of how performance changes with respect to $\lambda_{\text{cosntr}}$ and $\sigma_{\text{max}}$ for each constraint.
Appendix D Experiment details
D.1 Implementation
@@
For training the score model $s_{\theta}$ in the learned mirror space, we followed the implementation of Song et al. (2021a). We used the NCSN++ architecture with $64$ filters in the first layer and the VP SDE with $\beta_{\min}=0.1$ and $\beta_{\max}=20$. Training was done using the Adam optimizer with a learning rate of $0.0002$ and gradient norm clipping with a threshold of $1$.
@@
For $\mathbf{g}_{\phi}$, we followed the implementation of the gradient of a strongly-convex ICNN of Tan et al. (2023), configuring the ICNN to be $0.9$-strongly convex. Following the settings of CycleGAN (Zhu et al., 2017), $\mathbf{f}_{\psi}$ was implemented as a ResNet-based generator with $6$ residual blocks and $32$ filters in the last convolutional layer. For all constraints except the divergence-free constraint, we had the ResNet-based generator output the residual image (i.e., $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})=\text{ResNet}(\tilde{\mathbf{x}})+\tilde{\mathbf{x}}$). We found that for the divergence-free constraint, a non-residual-based inverse map (i.e., $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})=\text{ResNet}(\tilde{\mathbf{x}})$) achieves better constraint loss. The NAMM was trained using Adam optimizer with a learning rate of $0.001$ for the divergence-free constraint and a learning rate of $0.0002$ for all other constraints.
Table 3 shows the hyperparameter choices for each constraint. The regularization weight $\lambda_{\text{reg}}$ in the NAMM objective (Equation 1) was fixed at $0.001$. We used $3$ ICNN layers for images $64\times 64$ or smaller and $2$ ICNN layers for images $128\times 128$ or larger for the sake of efficiency. These hyperparameter values do not need to be heavily tuned, as we chose these settings through a coarse parameter search (e.g., trying $\lambda_{\text{constr}}=0.01$ or $\lambda=1$ to see which would lead to reasonable loss curves).
@@
The main results shown in Figure 3 were taken from the finetuned NAMM, ensuring that the total training time of the NAMM, MDM, and finetuning did not exceed the total training time of the baseline DM. While we kept track of the validation loss, this was not used to determine stopping time. We found that the NAMM training and MDM training were not prone to overfitting, so we chose the total number of epochs based on observing a reasonable level of convergence of the loss curves. We found that some overfitting is possible during finetuning but did not perform early stopping. All results were obtained from unseen test data because we fed random samples from the MDM into the inverse map and made sure not to use the same random seed as the one used to generate finetuning data. For all constraints, we generated $12800$ training examples from the MDM for finetuning.
Table 4 details the exact number of training epochs for each constraint. Figure 4 in the main results compares the constraint distances of our method without finetuning, our method with finetuning, and the baseline DM as a function of compute time.
@@
Total Brightness $30$ $200$ $1000$ $450$
1D Burgers’ $100$ $300$ $700$ $1500$
Divergence-free $100$ $700$ $500$ $2000$
Periodic $50$ $300$ $1000$ $1000$
Count $50$ $300$ $700$ $1500$
@@
For the comparison of parameterizing the mirror map as the gradient of an ICNN versus as a ResNet-based generator (Figure 7), we used a ResNet-based generator that outputs the residual image. This means that the inverse mirror map was parameterized as a residual-based network ($\mathbf{f}_{\psi}(\tilde{\mathbf{x}})=\text{ResNet}(\tilde{\mathbf{x}};\psi)+\tilde{\mathbf{x}}$), and so was the ResNet-based forward mirror map ($\mathbf{g}_{\phi}(\mathbf{x})=\text{ResNet}(\mathbf{x};\phi)+\mathbf{x}$).
D.2 Ablation of sequence of noisy mirror distributions
Recall that the NAMM training objective involves optimizing $\mathbf{f}_{\psi}$ over the sequence of noisy mirror distributions defined in Equation 2. Thus instead of considering a single noisy mirror distribution ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}\ast\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ with a fixed noise level $\sigma$, we perturb samples from ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$ with varying levels of Gaussian noise with variances ranging from 00 to $\sigma_{\text{max}}^{2}$. In Figure 10, we compare this choice, which involves integrating over $\sigma\in[0,\sigma_{\text{max}}]$, to the use of a fixed $\sigma=\sigma_{\text{max}}$.
Figure 10: Fixed $\sigma$ vs. integrating over $[0,\sigma_{\text{max}}]$. The NAMM objective involves optimizing $\mathbf{f}_{\psi}$ over the sequence of noisy mirror distributions defined in Eq. 2. We compare this approach of integrating over $\sigma\in[0,\sigma_{\text{max}}]$ to setting a fixed noise standard deviation $\sigma=\sigma_{\text{max}}$ in the context of the 1D Burgers’ constraint (here $\sigma_{\text{max}}=0.1$). Both before and after finetuning, the constraint distances $\ell_{\text{burgers}}$ of inverted MDM samples are smaller if the NAMM was trained with varying $\sigma$. The histograms show the constraint distances of $128$ samples from each method.
D.3 Data assimilation with mirror DPS
Following the original DPS (Chung et al., 2022), we use a hyperparameter $\zeta\in\mathbb{R}_{>0}$ to re-weight the time-dependent measurement likelihood. At diffusion time $t$, the measurement weight is given by